OpenStudy (anonymous):

What is the slope of the graph of 8x – 4y = 11? a.1/2 b.-1/2 c.2 d.-2

OpenStudy (anonymous):

answer is c. 2

OpenStudy (anonymous):

put the equation in y=mx+b m is the slope and b is the y - intercept

OpenStudy (anonymous):

but how did u get 2 that's what I wanna know

OpenStudy (anonymous):

Solve the equation for y

OpenStudy (anonymous):

\[8x - 4y = 11\] becomes \[y = 2x - 11\]

OpenStudy (mertsj):

1. Subtract 8x from both sides 2. Divide both sides by -4 3. The coefficient of x is the slope.

OpenStudy (anonymous):

dividing -8 by -4 = 2! slope=m, m=2

OpenStudy (anonymous):

(for y=mx+b form)

OpenStudy (mertsj):

@Kaveman You have to divide the 11 also!!!

OpenStudy (anonymous):

@Mertsj OH NO! The shame and dishonour I have brought upon my family! Nice catch, my bad. \[y = 2x - \frac{ 11 }{ 4 }\] -bows to the almighty Mertsj-